OpenStudy (anonymous):

What is the distance between points (9, 4) and (–3, 4) on a coordinate plane? units

OpenStudy (anonymous):

help

pooja195 (pooja195):

Use the distance formula \[\huge d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

OpenStudy (anonymous):

@Whitemonsterbunny17

OpenStudy (anonymous):

@super123

OpenStudy (anonymous):

help

pooja195 (pooja195):

\[\huge d=\sqrt{(-9-9)^2+(4-4)^2}\]

OpenStudy (anonymous):

how

pooja195 (pooja195):

I used the coordinates and plugged them into the formula

pooja195 (pooja195):

-9-9=?

OpenStudy (anonymous):

-18

OpenStudy (anonymous):

next

pooja195 (pooja195):

\[\huge d=\sqrt{(-18)^2+(4-4)^2}\] 4-4=?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

help

OpenStudy (anonymous):

@super123

OpenStudy (anonymous):

help

pooja195 (pooja195):

18^2=?

OpenStudy (anonymous):

324

pooja195 (pooja195):

What is the square root of 324?

OpenStudy (anonymous):

18

pooja195 (pooja195):

Good so the final answer is 18

OpenStudy (anonymous):

thanks

pooja195 (pooja195):

np